Stand-alone GNU Info

for version 7.1.1, 7 September 2024

Brian J. Fox and Texinfo maintainers

This manual is for Stand-alone GNU Info (version 7.1.1, 7 September 2024), a program for viewing documents in Info format (usually created from Texinfo source files). Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, 1993, 1996, 1997, 1998, 1999, 2001-2023 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License” in the Texinfo manual.

This document is part of a collection distributed under the GNU Free Documentation License. If you want to distribute this document separately from the collection, you can do so by adding a copy of the license to the document, as described in section 6 of the license.

Table of Contents

1 Stand-alone Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Selecting a Node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Further Commands for Selecting a Node . . . . . . 3 4 Moving the Cursor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5 Moving Text Within a Window. . . . . . . . . . . . . . . . . 6 6 Searching an Info File . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 7 Index Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 8 Selecting Cross-references. . . . . . . . . . . . . . . . . . . . . . 11

8.1 Parts of an Xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 8.2 Selecting Xrefs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

9 Manipulating Multiple Windows . . . . . . . . . . . . . . 13

9.1 The Mode Line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.2 Window Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 9.3 The Echo Area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

10 Printing Nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 11 Miscellaneous Commands. . . . . . . . . . . . . . . . . . . . . 18 12 Invoking Info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 13 Manipulating Variables . . . . . . . . . . . . . . . . . . . . . . . 27 14 Colors and Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 15 Customizing Key Bindings and Variables … 33

15.1 infokey format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

16 Alternate MS-DOS/Windows keybindings . . 37 Appendix A Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

1 Stand-alone Info

The Info program described here is a stand-alone program, part of the Texinfo distribution, which is used to view Info files on a text terminal. Info files are typically the result of processing Texinfo files with the program texi2any (also in the Texinfo distribution).

Texinfo itself (see Texinfo) is a documentation system that uses a single source file to produce both on-line information and printed output. You can typeset and print the files that you read in Info.

GNU Emacs also provides an Info reader (just type M-x info in Emacs). Emacs Info and stand-alone Info have nearly identical user interfaces, although customization and other details are different (this manual explains the stand-alone Info reader). The Emacs Info reader supports the X Window System and other such bitmapped interfaces, not just plain ASCII, so if you want a prettier display for Info files, you should try it. You can use Emacs Info without using Emacs for anything else. (Type C-x C-c to exit; this also works in the stand-alone Info reader.) See Info for a tutorial and more background information about the Info system, as well as information about the Info reader that is part of GNU Emacs,

Please report bugs in this stand-alone Info program to bug-texinfo@gnu.org. Bugs in the Emacs Info reader should be sent to bug-gnu-emacs@gnu.org.

2 Selecting a Node

The most basic node commands are ‘n’, ‘p’, and ‘u’. These move you around the node structure of the file:

n (next-node)

Select the ‘Next’ node.

p (prev-node)

Select the ‘Prev’ node.

u (up-node)

Select the ‘Up’ node.

The top line of each Info node contains pointers which describe where the next, previous, and up nodes are. The nodes in an Info file are conventionally arranged in a hierarchical structure; for example, the ‘Next’ pointer in a section of a chapter takes you to the next section in the same chapter, while the ‘Up’ pointer takes you to the higher-level node for the chapter as a whole.

You can select a node that you have already viewed by using the ‘l’ command—this stands for “last”:

l (history-node)

Move backwards through the history of visited nodes for this window. The current node is discarded from the history. This is useful when you follow a reference to another node to read about a related issue, and would like then to resume reading at the same place where you started.

Two additional commands, ‘t’ and ‘d’, select special nodes:

t (top-node)

Select the node ‘Top’ in the current Info file.

d (dir-node)

Select the directory node (i.e., the node ‘(dir)’). This contains a menu referencing all the available Info files that are installed on your system.

In the command descriptions in this manual, the M-x command names are displayed in parentheses. M-x is itself a command (execute-command) letting you run a command by name. For example, you could select the next node with M-x next-node.

C-x means press the Ctrl key and the key x. M-x means press the Meta key and the key x. (Meta is usually labeled as Alt). SPC is the space bar. The other keys are usually called by the names printed on them. This notation to describe keystrokes is the same as that used within the Emacs manual (see Section “User Input” in The GNU Emacs Manual).

Many Info commands can take a numeric argument; see Chapter 11 [Miscellaneous Commands], page 18, to find out how to supply one.

3 Further Commands for Selecting a Node

This chapter describes some other commands which select a different node.

< (first-node)

Selects the first node which appears in this file. This node is most often ‘Top’, but it does not have to be. With a numeric argument N, select the Nth node (the first node is node 1). An argument of zero is the same as the argument of 1.

> (last-node)

Select the last node which appears in this file. With a numeric argument N, select the Nth node (the first node is node 1). An argument of zero is the same as no argument, i.e., it selects the last node.

] (global-next-node)

Move forward through the node structure. If the node that you are currently viewing has a menu, select the first menu item. Otherwise, if this node has a ‘Next’ pointer, follow it. If there is no menu and no ‘Next’ pointer, then follow ‘Up’ pointers until there is a ‘Next’ pointer, and then follow it.

[ (global-prev-node)

Move backward through the node structure. If the node that you are currently viewing has a ‘Prev’ pointer, that node is selected. Otherwise, if the node has an ‘Up’ pointer, that node is selected, and if it has a menu, the last item in the menu is selected.

You can get the same behavior as global-next-node and global-prev-node

while simply scrolling through the file with SPC and DEL (see [scroll-behavior], page 29).

g (goto-node)

Read the name of a node and select it. If the desired node resides in some other file, you must type the node as it appears in that Info file, and include the name of the other file. For example,

     g(emacs)Buffers

finds the node ‘Buffers’ in the Info file emacs. While reading the node name, completion (see Section 9.3 [The Echo Area], page 14) is only done for the nodes which reside in one of the Info files that were loaded in the current Info session.

O (goto-invocation)

Read the name of a program and look for a node in the current Info file which describes the invocation and the command-line options for that program. The default program name is derived from the name of the current Info file. This command does the same as the ‘--show-options’ command-line option (see [–show-options], page 25), but it also allows to specify the program name; this is important for those manuals which describe several programs. If you need to find the Invocation node of a program that is documented in another Info file, you need to visit that file before invoking ‘O’. For example, if you are reading the Emacs manual and want to see the command-line options

of the texi2any program, type g (texinfo) RET and then O texi2any RET.

If you don’t know what Info file documents the command, or if invoking ‘O’ doesn’t display the right node, go to the ‘(dir)’ node (using the ‘d’ command) and invoke ‘O’ from there.

G (menu-sequence)

Read a sequence of menu entries and follow it. Info prompts for a sequence of menu items separated by commas. (Since commas are not allowed in a node name, they are a natural choice for a delimiter in a list of menu items.) Info then looks up the first item in the menu of the node ‘(dir)’ (if the ‘(dir)’ node cannot be found, Info uses ‘Top’). If such an entry is found, Info goes to the node it points to and looks up the second item in the menu of that node, etc. In other words, you can specify a complete path which descends through the menu hierarchy of a particular Info file starting at the ‘(dir)’ node. This has the same effect as if you typed the menu item sequence on Info’s command line, see [Info command-line arguments processing], page 21. For example,

      G Texinfo,Overview,Reporting Bugs RET

displays the node ‘Reporting Bugs’ in the Texinfo manual. (You don’t actually need to type the menu items in their full length, or in their exact letter-case. However, if you do type the menu items exactly, Info will find it faster.) If any of the menu items you type are not found, Info stops at the last entry it did find and reports an error.

C-x C-f (view-file)

Read the name of a file and selects the entire file. The command

     C-x C-f filename

is equivalent to typing

     g(filename)*
C-x C-b (list-visited-nodes)

Make a window containing a menu of all of the currently visited nodes. This window becomes the selected window, and you may use the standard Info commands within it.

C-x b (select-visited-node)

Select a node which has been previously visited in a visible window. This is similar to ‘C-x C-b’ followed by ‘m’, but no window is created.

M-x man Read the name of a man page to load and display. This uses the man command on your system to retrieve the contents of the requested man page. See also [–raw-escapes], page 24.

4 Moving the Cursor

GNU Info has several commands which allow you to move the cursor about the screen.

With a numeric argument, the motion commands are simply executed that many times; for example, a numeric argument of 4 given to next-line causes the cursor to move down 4 lines. With a negative numeric argument, the motion is reversed; an argument of −4 given to the next-line command would cause the cursor to move up 4 lines.

C-n (next-line)

DOWN (an arrow key)

Move the cursor down to the next line.

C-p (prev-line)

UP (an arrow key)

Move the cursor up to the previous line.

C-a (beginning-of-line)

Move the cursor to the start of the current line.

C-e (end-of-line)

Move the cursor to the end of the current line.

C-f (forward-char)

RIGHT (an arrow key)

Move the cursor forward a character.

C-b (backward-char)

LEFT (an arrow key)

Move the cursor backward a character.

M-f (forward-word)

Move the cursor forward a word.

M-b (backward-word)

Move the cursor backward a word.

M-< (beginning-of-node)

b Move the cursor to the start of the current node.

M-> (end-of-node)

e Move the cursor to the end of the current node.

M-r (move-to-window-line)

Move the cursor to a specific line of the window. Without a numeric argument, M-r moves the cursor to the start of the line in the center of the window. With a numeric argument of n, M-r moves the cursor to the start of the nth line in the window.

5 Moving Text Within a Window

Sometimes you are looking at a screenful of text, and only part of the current paragraph you are reading is visible on the screen. The commands detailed in this section are used to shift which part of the current node is visible on the screen.

SPC (scroll-forward)

NEXT Shift the text in this window up. That is, show more of the node which is currently below the bottom of the window. With a numeric argument, show that many more lines at the bottom of the window; a numeric argument of 4 would shift all of the text in the window up 4 lines (discarding the top 4 lines), and show you four new lines at the bottom of the window. Without a numeric argument, SPC takes the bottom two lines of the window and places them at the top of the window, redisplaying almost a completely new screenful of lines. If you are at the end of a node, SPC takes you to the “next” node, so that you can read an entire manual from start to finish by repeating SPC. The NEXT key is known as the PageDown key on some keyboards.

C-v (scroll-forward-page-only)

Shift the text in this window up. This is identical to the SPC operation above, except that it never scrolls beyond the end of the current node.

M-x scroll-forward-page-only-set-window

Scroll forward, like with C-v, but if a numeric argument is specified, it becomes the default scroll size for subsequent scroll-forward and scroll-backward commands and their ilk.

DEL (scroll-backward)

PREVIOUS Shift the text in this window down. The inverse of scroll-forward. If you are at the start of a node, DEL takes you to the “previous” node, so that you can read an entire manual from finish to start by repeating DEL. The default scroll size

can be changed by invoking the (scroll-backward-page-only-set-window)

command with a numeric argument. If your keyboard lacks the DEL key, look for a key called BS, or ‘Backspace’, sometimes designated with an arrow which points to the left, which should perform the same function. The PREVIOUS key is the PageUp key on many keyboards. Emacs refers to it by the name PRIOR.

M-v (scroll-backward-page-only)

Shift the text in this window down. The inverse of scroll-forward-page-only. Does not scroll beyond the start of the current node. The default scroll

size can be changed by invoking the scroll-backward-page-only-set-window

command with a numeric argument.

M-x scroll-backward-page-only-set-window

Scroll backward, like with M-v, but if a numeric argument is specified, it becomes the default scroll size for subsequent scroll-forward and scroll-backward commands.

M-x down-line

Scroll forward by one line. With a numeric argument, scroll forward that many lines.

M-x up-line

Scroll backward one line. With a numeric argument, scroll backward that many lines.

M-x scroll-half-screen-down

Scroll forward by half of the screen size. With a numeric argument, scroll that many lines. If an argument is specified, it becomes the new default number

of lines to scroll for subsequent scroll-half-screen-down and scroll-half-
screen-up commands.
M-x scroll-half-screen-up

Scroll back by half of the screen size. With a numeric argument, scroll that many lines. If an argument is specified, it becomes the new default number

of lines to scroll for subsequent scroll-half-screen-down and scroll-half-
screen-up commands.

The scroll-forward and scroll-backward commands can also move forward and backward through the node structure of the file. If you press SPC while viewing the end of a node, or DEL while viewing the beginning of a node, what happens is controlled by the variable scroll-behavior (see [scroll-behavior], page 29).

The scroll-forward-page-only and scroll-backward-page-only commands never

scroll beyond the current node.

6 Searching an Info File

GNU Info allows you to search for a sequence of characters throughout an entire Info file. Here are the commands to do this:

s (search)

/ Read a string in the echo area and search for it, either as a regular expression (by default) or a literal string. If the string includes upper-case characters, the Info file is searched case-sensitively; otherwise Info ignores the letter case. With a numeric argument of N, search for Nth occurrence of the string. Negative arguments search backwards.

? (search-backward)

Read a string in the echo area and search backward through the Info file for that string. If the string includes upper-case characters, the Info file is searched casesensitively; otherwise Info ignores the letter case. With a numeric argument of N, search for Nth occurrence of the string. Negative arguments search forward.

C-x n (search-next)

} Search forwards for the string used for the last search command. Case sensitivity and use of regular expressions are kept the same. With a numeric argument of n, search for nth next occurrence. By default, the search starts at the position immediately following the cursor. However, if the variable search-skip-screen (see Chapter 13 [search-skip-screen], page 27) is set, it starts at the beginning of the next page, thereby skipping all visibly displayed lines.

C-x N (search-previous)

{ Just like search-next, but in reverse. You can use search-next and searchprevious together to move forward and backward through matches. searchprevious usually goes to the place in the file that was displayed before an immediately preceding search-next, and vice versa.^{1}

R (toggle-regexp)

Toggle between using regular expressions and literal strings for searching. Info uses so-called ‘extended’ regular expression syntax (see Section “Regular Expressions” in GNU Grep).

S (search-case-sensitively)

Read a string in the echo area and search for it case-sensitively, even if the string includes only lower-case letters. With a numeric argument of N, search for Nth occurrence of the string. Negative arguments search backwards.

C-s (isearch-forward)

Interactively search forward through the Info file for a string as you type it. If the string includes upper-case characters, the search is case-sensitive; otherwise Info ignores the letter case.

1 This sometimes doesn’t happen when search-skip-screen is On, and the search goes across nodes.

C-r (isearch-backward)

Interactively search backward through the Info file for a string as you type it. If the string includes upper-case characters, the search is case-sensitive; otherwise Info ignores the letter case.

M-/ (tree-search)

Recursively search this node and any subnodes listed in menus for a string.

M-} (tree-search-next)
M-{ (tree-search-previous)

Go forwards and backwards through the matches for an active tree search.

The most basic searching command is ‘s’ or ‘/’ (search). The ‘s’ command prompts you for a string in the echo area, and then searches the remainder of the Info file for an occurrence of that string. If the string is found, the node containing it is selected, and the cursor is left positioned at the start of the found string. Subsequent ‘s’ commands show you the default search string; pressing RET instead of typing a new string will use the default search string.

Incremental searching is similar to basic searching, but the string is looked up while you are typing it, instead of waiting until the entire search string has been specified.

The tree search can be used from the dir node to search through all Info files installed on the system. It can also be used to search through a particular chapter of a manual when you are not interested in matches in other chapters.

If the highlight-searches variable is set, matches from search commands will be highlighted. See Chapter 13 [highlight-searches], page 27. Use the M-x clear-search command to clear any search highlights.

Both incremental and non-incremental search by default ignore the case of letters when comparing the Info file text with the search string. However, an uppercase letter in the search string makes the search case-sensitive. You can force a case-sensitive non-incremental search, even for a string that includes only lower-case letters, by using the ‘S’ command (search-case-sensitively). The ‘n’ and ‘N’ commands operate case-sensitively if the last search command was ‘S’.

Normally, the search pattern should not be shorter than some predefined limit. By default, this limit is set to 1 character. See [min-search-length], page 29, for more information on this.

7 Index Commands

GNU Info has commands to search through the indices of an Info file, which helps you find areas within an Info file which discuss a particular topic.

i (index-search)

Look up a string in the indices for this Info file, and select a node to which the found index entry points.

I (virtual-index)

Look up a string in the indices for this Info file, and show all the matches in a new virtual node, synthesized on the fly.

, (next-index-match)

Move to the node containing the next matching index item from the last ‘i’ command.

M-x index-apropos

Grovel the indices of all the known Info files on your system for a string, and build a menu of the possible matches.

The most efficient means of finding something quickly in a manual is the ‘i’ command (index-search). This command prompts for a string, and then looks for that string in all the indices of the current Info manual. If it finds a matching index entry, it displays the node to which that entry refers and prints the full text of the entry in the echo area. You can press ‘,’ (next-index-match) to find more matches. A good Info manual has all of its important concepts indexed, so the ‘i’ command lets you use a manual as a reference.

If you don’t know what manual documents something, try the M-x index-apropos command. It prompts for a string and then looks up that string in all the indices of all the Info documents installed on your system. It can also be invoked from the command line; see [–apropos], page 22.

8 Selecting Cross-references

We have already discussed the ‘Next’, ‘Prev’, and ‘Up’ pointers which appear at the top of a node, referring you to a different node.

In addition, a node may contain a menu, as well as cross-references (xrefs for short) interspersed through the text of the node. Cross-references may possibly refer to a node in another Info file.

8.1 Parts of an Xref

Here is a sample menu entry showing the parts of a cross-reference:

     * Foo Label: Foo Target.        More information about Foo.

The reference has two parts: the first part is called the label; it is the name that you can use to refer to the cross-reference, and the second is the target; it is the full name of the node that the cross-reference points to. The target is separated from the label by a single colon ‘:’; first the label appears, and then the target.

The ‘.’ ends the name of the target. The ‘.’ is not part of the target; it serves only to let Info know where the target name ends.

A shorthand way of specifying references allows two adjacent colons to stand for a target name which is the same as the label name:

     * Foo Commands::                Commands pertaining to Foo.

In the above example, the name of the target is the same as the name of the label, in

this case Foo Commands.

You will normally see two types of cross-reference while viewing nodes: menu references, and note references. Menu references appear within a node’s menu; they begin with a ‘*’ at the beginning of a line, and continue with a label, a target, and a comment which describes what the contents of the node pointed to contains.

Note references appear within the body of the node text; they begin with *Note, and continue with a label and a target.

Like ‘Next’, ‘Prev’, and ‘Up’ pointers, cross-references can point to any valid node. They are used to refer you to a place where more detailed information can be found on a particular subject. Here is a cross-reference which points to a node within the Texinfo documentation: See Section “Cross-refrences” in the Texinfo Manual, for more information on creating your own Texinfo cross-references.

8.2 Selecting Xrefs

The following table lists the Info commands which operate on menu items.

1 (menu-digit)

2 … 9 M-1, vi-like operation M-2 … M-9, vi-like operation

Within an Info window, pressing a single digit, (such as ‘1’), selects that menu item, and places its node in the current window. For convenience, there is one exception; pressing ‘0’ selects the last item in the node’s menu. When ‘--vi-keys’ is in effect, digits set the numeric argument, so these commands are remapped to their ‘M-’ varieties. For example, to select the last menu item, press M-0.

0 (last-menu-item)

M-0, vi-like operation Select the last item in the current node’s menu.

m (menu-item)

Reads the name of a menu item in the echo area and selects its node. Completion is available while reading the menu label. See Section 9.3 [The Echo Area], page 14.

M-x find-menu

Move the cursor to the start of this node’s menu.

This table lists the Info commands which operate on cross-references.

f (xref-item)

r Reads the name of a note cross-reference in the echo area and selects its node. Completion is available while reading the cross-reference label. See Section 9.3 [The Echo Area], page 14.

Finally, the next few commands operate on menu or note references alike:

TAB (move-to-next-xref)

Move the cursor to the start of the next nearest menu item or note reference in this node. You can then use RET (select-reference-this-line) to select the menu or note reference.

M-TAB (move-to-prev-xref)

BackTab Move the cursor the start of the nearest previous menu item or note reference in this node. The BackTab key can be produced on some terminals with Shift-TAB.

RET (select-reference-this-line)

Select the menu item or note reference appearing on this line.

9 Manipulating Multiple Windows

A window is a place to show the text of a node. Windows have a view area where the text of the node is displayed, and an associated mode line, which briefly describes the node being viewed.

GNU Info supports multiple windows appearing in a single screen; each window is separated from the next by its mode line. At any time, there is only one active window, that is, the window in which the cursor appears. There are commands available for creating windows, changing the size of windows, selecting which window is active, and for deleting windows.

9.1 The Mode Line

A mode line is a line of inverse video which appears at the bottom of an Info window. It describes the contents of the window just above it; this information includes the name of the file and node appearing in that window, the number of screen lines it takes to display the node, and the percentage of text that is above the top of the window.

Here is a sample mode line for a window containing a file named dir, showing the node

‘Top’.
     -----Info: (dir)Top, 40 lines --Top-------------------------------------
                 ^^^^^^^^
               (file)Node #lines    where

Truncation of long lines (as opposed to wrapping them to the next display line, see Chapter 5 [Scrolling Commands], page 6) is indicated by a ‘$’ at the left edge of the mode line:

     --$--Info: (texinfo)Top, 480 lines --Top--------------------------------

When Info makes a node internally, such that there is no corresponding info file on disk, the name of the node is surrounded by asterisks (‘*’). The name itself tells you what the contents of the window are; the sample mode line below shows an internally constructed node showing possible completions:

     -----Info: *Completions*, 7 lines --All---------------------------------

9.2 Window Commands

It can be convenient to view more than one node at a time. To allow this, Info can display more than one window. Each window has its own mode line (see Section 9.1 [The Mode Line], page 13) and history of nodes viewed in that window (see Chapter 2 [history-node], page 2).

C-x o (next-window)

Select the next window on the screen. Note that the echo area can only be selected if it is already in use, and you have left it temporarily. Normally, ‘C-x o’ simply moves the cursor into the next window on the screen, or if you are already within the last window, into the first window on the screen. Given a numeric argument, ‘C-x o’ moves over that many windows. A negative argument causes ‘C-x o’ to select the previous window on the screen.

M-x prev-window

Select the previous window on the screen. This is identical to ‘C-x o’ with a negative argument.

C-x 2 (split-window)

Split the current window into two windows, both showing the same node. Each window is one half the size of the original window, and the cursor remains in the original window. The variable automatic-tiling can cause all of the windows on the screen to be resized for you automatically (see Chapter 13

[automatic-tiling], page 27).
C-x 0 (delete-window)

Delete the current window from the screen. If you have made too many windows and your screen appears cluttered, this is the way to get rid of some of them.

C-x 1 (keep-one-window)

Delete all of the windows excepting the current one.

ESC C-v (scroll-other-window)

Scroll the other window, in the same fashion that ‘C-v’ might scroll the current window. Given a negative argument, scroll the “other” window backward.

C-x ^(grow-window)

Grow (or shrink) the current window. Given a numeric argument, grow the current window that many lines; with a negative numeric argument, shrink the window instead.

C-x t (tile-windows)

Divide the available screen space among all of the visible windows. Each window is given an equal portion of the screen in which to display its contents. The variable automatic-tiling can cause tile-windows to be called when a window is created or deleted. See Chapter 13 [automatic-tiling], page 27.

9.3 The Echo Area

The echo area is a one line window which appears at the bottom of the screen. It is used to display informative or error messages, and to read lines of input from you when that is necessary. Almost all of the commands available in the echo area are identical to their Emacs counterparts, so please refer to that documentation for greater depth of discussion on the concepts of editing a line of text. The following table briefly lists the commands that are available while input is being read in the echo area:

C-f (echo-area-forward)

RIGHT (an arrow key)

Move forward a character.

C-b (echo-area-backward)

LEFT (an arrow key)

Move backward a character.

C-a (echo-area-beg-of-line)

Move to the start of the input line.

C-e (echo-area-end-of-line)

Move to the end of the input line.

M-f (echo-area-forward-word)

Move forward a word.

M-b (echo-area-backward-word)

Move backward a word.

C-d (echo-area-delete)

Delete the character under the cursor.

DEL (echo-area-rubout)

Delete the character behind the cursor. On some keyboards, this key is designated BS, for ‘Backspace’. Those keyboards will usually bind DEL in the echo area to echo-area-delete.

C-g (echo-area-abort)

Cancel or quit the current operation. If completion is being read, this command discards the text of the input line which does not match any completion. If the input line is empty, it aborts the calling function.

RET (echo-area-newline)

Accept (or forces completion of) the current input line.

C-q (echo-area-quoted-insert)

Insert the next character verbatim. This is how you can insert control characters into a search string.

M-TAB (echo-area-tab-insert)

Insert a TAB character.

C-t (echo-area-transpose-chars)

Transpose the characters at the cursor.

printing character Insert the character.

The next group of commands deal with killing, and yanking text. (Sometimes these operations are called cut and paste, respectively.) For an in-depth discussion, see Section “Killing and Deleting” in the GNU Emacs Manual.

M-d (echo-area-kill-word)

Kill the word following the cursor.

M-DEL (echo-area-backward-kill-word)

M-BS Kill the word preceding the cursor.

On some keyboards, the ‘Backspace’ key is used instead of DEL, so M-Backspace has the same effect as M-DEL.

C-k (echo-area-kill-line)

Kill the text from the cursor to the end of the line.

C-x DEL (echo-area-backward-kill-line)

Kill the text from the cursor to the beginning of the line.

C-y (echo-area-yank)

Yank back the contents of the last kill.

M-y (echo-area-yank-pop)

Yank back a previous kill, removing the last yanked text first.

Sometimes when reading input in the echo area, the command that needed input will only accept one of a list of several choices. The choices represent the possible completions, and you must respond with one of them. Since there are a limited number of responses you can make, Info allows you to abbreviate what you type, only typing as much of the response as is necessary to uniquely identify it. In addition, you can request Info to fill in as much of the response as is possible; this is called completion.

The following commands are available when completing in the echo area:

TAB (echo-area-complete)

Insert as much of a completion as is possible. Otherwise, display a window containing a list of the possible completions of what you have typed so far. For example, if the available choices are:

     bar foliate food forget

and you have typed an ‘f’, followed by TAB, this would result in ‘fo’ appearing in the echo area, since all of the choices which begin with ‘f’ continue with ‘o’. Now if you type TAB again, Info will pop up a window showing a node called ‘*Completions*’ which lists the possible completions like this:

     3 completions:
     foliate         food
     forget

i.e., all of the choices which begin with ‘fo’. Now, typing ‘l’ followed by ‘TAB’ results in ‘foliate’ appearing in the echo area, since that is the only choice which begins with ‘fol’.

ESC C-v (echo-area-scroll-completions-window)

Scroll the completions window, if that is visible, or the “other” window if not.

10 Printing Nodes

In general, we recommend that you use TEX to format the document and print sections of it, by running tex on the Texinfo source file. However, you may wish to print out the contents of a node as a quick reference document for later use, or if you don’t have TEX installed. Info provides you with a command for doing this.

M-x print-node

Pipe the contents of the current node through the command in the environment variable INFO_PRINT_COMMAND. If the variable does not exist, the node is simply piped to lpr (on DOS/Windows, the default is to print the node to the local printer device, PRN). The value of INFO_PRINT_COMMAND may begin with the ‘>’ character, as in ‘>/dev/printer’, in which case Info treats the rest as the name of a file or a device. Instead of piping to a command, Info opens the file, writes the node contents, and closes the file, under the assumption that text written to that file will be printed by the underlying OS.

11 Miscellaneous Commands

GNU Info contains several commands which self-document GNU Info:

M-x describe-command

Read the name of an Info command in the echo area and then display a brief description of what that command does.

M-x describe-key

Read a key sequence in the echo area, and then display the name and documentation of the Info command that the key sequence invokes.

M-x describe-variable

Read the name of a variable in the echo area and then display a brief description of what the variable affects.

M-x where-is

Read the name of an Info command in the echo area, and then display a key sequence which can be typed in order to invoke that command.

H (get-help-window)

Create (or Move into) the window displaying *Help*, and place a node containing a quick reference card into it. This window displays the most concise information about GNU Info available.

h (get-info-help-node)

Try hard to visit the node (info)Help. The Info file info.texi distributed with GNU Emacs contains this node. Of course, the file must first be processed with texi2any, and then placed into the location of your Info directory.

= (display-file-info)

Show information about what’s currently being viewed in the echo area: the Info file name, and current line number and percentage within the current node.

M-x info-version

Display the name and version of the currently running Info program.

Here are the commands for creating a numeric argument:

C-u (universal-argument)

Start (or multiply by 4) the current numeric argument. ‘C-u’ is a good way to give a small numeric argument to cursor movement or scrolling commands; ‘C-u C-v’ scrolls the screen 4 lines, while ‘C-u C-u C-n’ moves the cursor down 16 lines. ‘C-u’ followed by digit keys sets the numeric argument to the number thus typed: C-u 1 2 0 sets the argument to 120.

M-1 (add-digit-to-numeric-arg)

1, vi-like operation

M-2 . . . M-9

2 … 9, vi-like operation

M-0

0, vi-like operation Add the digit value of the invoking key to the current numeric argument. Once Info is reading a numeric argument, you may just type the digits of the argument, without the Meta prefix. For example, you might give ‘C-l’ a numeric argument of 32 by typing:

     C-u 3 2 C-l

or

     M-3 2 C-l
M-- (add-digit-to-numeric-arg)

- To make a negative argument, type -. Typing - alone makes a negative argument with a value of −1. If you continue to type digit or Meta-digit keys after -, the result is a negative number produced by those digits.

- doesn’t work when you type in the echo area, because you need to be able to insert the ‘-’ character itself; use M-- instead, if you need to specify negative arguments in the echo area.

C-g is used to abort the reading of a multi-character key sequence, to cancel lengthy operations (such as multi-file searches) and to cancel reading input in the echo area.

C-g (abort-key)

Cancel current operation.

The ‘q’ command of Info simply quits running Info.

q (quit)

C-x C-c Exit GNU Info.

Here are commands affecting the display of nodes:

C-l (redraw-display)

Redraw the display from scratch, or shift the line containing the cursor to a specified location. With no numeric argument, ‘C-l’ clears the screen, and then redraws its entire contents. Given a numeric argument of n, the line containing the cursor is shifted so that it is on the nth line of the window.

M-x set-screen-height

Read a height value in the echo area and set the height of the displayed screen to that value. For example, if the operating system tells GNU Info that the screen is 60 lines tall, and it is actually only 40 lines tall, you could use this command to tell Info that the operating system is incorrect.

C-x w (toggle-wrap)

Toggles the state of line wrapping in the current window. Normally, lines which are longer than the screen width wrap, i.e., they are continued on the next line. Lines which wrap have a ‘\’ appearing in the rightmost column of the screen. You can cause such lines to be terminated at the rightmost column by changing the state of line wrapping in the window with C-x w. When a line which needs more space than one screen width to display is displayed, a ‘$’ appears in the rightmost column of the screen, and the remainder of the line is invisible. When long lines are truncated, the mode line displays the ‘$’ character near its left edge.

On MS-DOS/MS-Windows, this command actually tries to change the dimensions of the visible screen to the value you type in the echo area.

Finally, Info provides a way to display footnotes which might be associated with the current node that you are viewing:

ESC C-f (show-footnotes)

Show the footnotes (if any) associated with the current node in another window. You can have Info automatically display the footnotes associated with a node when the node is selected by setting the variable automatic-footnotes. See

Chapter 13 [automatic-footnotes], page 27.

12 Invoking Info

GNU Info accepts several options to control the initial node or nodes being viewed, and to specify which directories to search for Info files. Here is a template showing an invocation of GNU Info from the shell:

     info [option...] [manual] [menu-or-index-item...]

Info will look for an entry called manual in the directory files, which are named dir, that it finds in its search path. The search is case-insensitive and considers substrings. (If manual is not given, by default Info displays a composite directory listing, constructed by combining the dir files.) A basic example:

     info coreutils

This looks for an entry labelled coreutils, or Coreutils, etc., and if found, displays the referenced file (e.g., coreutils.info) at the location given. info coreu will find it too, if there is no better match.

Another example:

     info ls

Assuming the normal dir entry for ls, this will show the ls documentation, which happens to be within the coreutils manual rather than a separate manual. The dir entries can point to any node within a manual, so that users don’t have to be concerned with the exact structure used by different authors.

If no entry is found in the directories, Info looks for a file called manual in its search path. If not found, Info looks for it with the file extensions .info, -info, and .inf. For each of these known extensions, if a regular file is not found, Info looks for a compressed file.^{1}

You can specify the name of a node to visit with the --node or -n option. Alternatively, you can specify the file and node together using the same format that occurs in Info crossreferences. These two examples both load the ‘Files’ node within the ‘emacs’ manual:

     info emacs -n Files
     info '(emacs)Files'

If you want to load a file without looking in the search path, specify manual either as an absolute path, or as a path relative to the current directory which contains at least one slash character. (You can also use the --file option for similar behavior, described below.) Examples:

     info /usr/local/share/info/bash.info
     info ./document.info

Info looks for manual only in the explicitly specified directory, and adds that directory to its search path.

1 Info supports files compressed with gzip, xz, bzip2, lzip, lzma, compress and yabba programs, assumed

to have extensions .z, .gz, .xz, .bz2, .lz, .lzma, .Z, and .Y respectively.

On MS-DOS, Info allows for the Info extension, such as .inf, and the short compressed file extensions,

such as .z and .gz, to be merged into a single extension, since DOS doesn’t allow more than a single

dot in the basename of a file. Thus, on MS-DOS, if Info looks for bison, file names like bison.igz and

bison.inz will be found and decompressed by gunzip.

Info treats any remaining arguments as the names of menu items, or (see below) index entries. The first argument is a menu item in the ‘Top’ node of the file loaded, the second argument is a menu item in the first argument’s node, etc. You can move to the node of your choice by specifying the menu names which describe the path to that node. For example,

     info emacs buffers
     info texinfo Overview 'Using Texinfo'

The first example selects the menu item ‘Buffers’ in the node ‘(emacs)Top’. The second example loads the texinfo file and looks in its top-level menu for a ‘Overview’ item, looks in the menu of the node referenced, and finally displays the node referenced by the ‘Using

Texinfo’ item.

If there was only one menu-or-index-item argument and it wasn’t found as a menu item, Info looks for it as an index entry. For example:

     info libc printf

This loads the libc Info manual and first looks for printf in the top-level menu as usual; since it isn’t there (at this writing), it then looks in the indices. If it’s found there (which it is), the relevant node at the given location is displayed.

If Info is invoked when its standard output is not a terminal, it does not attempt to start an interactive session; rather, it writes the contents of the loaded nodes and subnodes to standard output, as if the --output=- and --subnodes options were given. This can be used to pipe the contents of Info nodes to another program, such as a pager.

A complete list of options follows.

--all

-a Find all files matching manual. Three usage patterns are supported, as follows. First, if --all is used together with --where, info prints the names of all matching files found on standard output (including ‘*manpages*’ if relevant) and exits. Second, if --all is used together with --output, the contents of all matched files are dumped to the specified output file. Otherwise, an interactive session is initiated. If more than one file matches, a menu node is displayed listing the matches and allowing you to select one. This menu node can be brought back at any time by pressing C-x f. If there is only one match, info starts as usual. When used with the --index-search option, info displays a menu of matching index entries (just as the virtual-index command does; see Chapter 7 [Index Commands], page 10). The --node option cannot be used together with this option.

--apropos=string

-k string Specify a string to search in every index of every Info file installed on your system. Info looks up the named string in all the indices it can find, prints the results to standard output, and then exits. If you are not sure which Info file explains certain issues, this option is your friend. (If your system has a lot of Info files installed, searching all of them might take some time!)

You can invoke the apropos command from inside Info; see Chapter 6 [Searching Commands], page 8.

--debug=number

-x number Print additional debugging information. The argument specifies the verbosity level, so a higher level includes all the information from lower levels. For all available debugging output, use -x -1. Info version 7.1.1 has these levels:

1 Print information about file handling, such as looking for dir files and nodes written with ‘--output’.

2 Print operations relating to INFOPATH.

3 Print information about node searching.

Debugging output goes to standard error.

--directory directory-path
-d directory-path

Add directory-path to the list of directory paths searched when Info needs to find a file. You may issue --directory multiple times; once for each directory which contains Info files, or with a list of such directories separated by a colon (or semicolon on MS-DOS/MS-Windows). Directories specified in the environment variable INFOPATH are added to the directories specified with --directory, if any. The value of INFOPATH is a list of directories usually separated by a colon; on MS-DOS/MS-Windows systems, the semicolon is used. If the value of INFOPATH ends with a colon (or semicolon on MS-DOS/MS-Windows), the initial list of directories is constructed by appending the build-time default to the value of INFOPATH. If you do not define INFOPATH, Info uses a default path defined when Info was built as the initial list of directories. Regardless of whether INFOPATH is defined, the default documentation directory defined when Info was built is added to the search path. If you do not want this directory to be included, set the infopath-no-defaults variable to On (see [infopath-no-defaults], page 28). If the list of directories contains the element PATH, that element is replaced by a list of directories derived from the value of the environment variable PATH. Each path element of the form dir/base is replaced by dir/share/info or dir/info, provided that directory exists.

--dribble=file

Specify a file where all user keystrokes will be recorded. This file can be used later to replay the same sequence of commands, see the ‘--restore’ option below.

--file manual

-f manual Specify a particular manual to visit without looking its name up in any dir files. With this option, it starts by trying to visit (manual)Top, i.e., the Top node in (typically) manual.info. As above, it tries various file extensions to find the file. If no such file (or node) can be found, Info exits without doing anything. As with the dir lookup described above, any extra menu-or-index-item arguments are used to locate a node within the loaded file. If manual is an absolute file name, or begins with ./ or ../, or contains an intermediate directory, Info will only look for the file in the directory specified, and add this directory to INFOPATH. (This is the same as what happens when --file is not given.)

--help

-h Output a brief description of the available Info command-line options.

--index-search string

After processing all command-line arguments, go to the index in the selected Info file and search for index entries which match string. If such an entry is found, the Info session begins with displaying the node pointed to by the first matching index entry; press , to step through the rest of the matching entries. If no such entry exists, print ‘no entries found’ and exit with nonzero status. This can be used from another program as a way to provide online help, or as a quick way of starting to read an Info file at a certain node when you don’t know the exact name of that node. When used with the --all option, info displays a menu of matching index entries (just as the virtual-index command does; see Chapter 7 [Index Commands], page 10). This command can also be invoked from inside Info; see Chapter 6 [Searching Commands], page 8.

--init-file INIT-FILE

Read key bindings and variable settings from INIT-FILE instead of the .infokey file in your home directory. See Chapter 15 [Custom Key Bindings], page 33.

--node nodename
-n nodename

Specify a particular node to visit in the initial file that Info loads. You may specify --node multiple times: for an interactive Info, each nodename is visited in its own window; for a non-interactive Info (such as when --output is given) each nodename is processed sequentially. You can specify both the file and node to the --node option using the usual Info syntax, but don’t forget to escape the open and close parentheses and whitespace from the shell; for example:

info --node "(emacs)Buffers"
--output file

-o file Direct output to file. Each node that Info visits will be output to file instead of interactively viewed. A value of - for file means standard output.

--no-raw-escapes
--raw-escapes, -R

By default, Info passes SGR terminal control sequences (also known as ANSI escape sequences) found in documents directly through to the terminal. If you use the --no-raw-escapes options, these sequences are displayed as other control characters are; for example, an ESC byte is displayed as ‘^[’. The --raw-escapes and -R options do not do anything, but are included for completeness.

--restore=dribble-file

Read keystrokes from dribble-file, presumably recorded during previous Info session (see the description of the ‘--dribble’ option above). When the keystrokes in the files are all read, Info reverts its input to the usual interactive operation.

--show-malformed-multibytes
--no-show-malformed-multibytes

Show malformed multibyte sequences in the output. By default, such sequences are dropped.

--show-options
--usage

-O Look for the node that describes how to invoke the program. The name of the program is taken from the other non-option arguments on the command line.

For example, ‘info emacs -O’ loads the Emacs Invocation node of the emacs

manual. This option is provided to make it easier to find the most important usage information in a manual without navigating through menu hierarchies. The effect is similar to the M-x goto-invocation command (see [goto-invocation], page 3) from inside Info.

--speech-friendly

-b On MS-DOS/MS-Windows only, this option causes Info to use standard file I/O functions for screen writes. (By default, Info uses direct writes to the video memory on these systems, for faster operation and colored display support.) This allows the speech synthesizers used by blind persons to catch the output and convert it to audible speech.

--strict-node-location

This option causes Info not to search “nearby” to locate nodes, and instead strictly use the information provided in the Info file. The practical use for this option is for debugging programs that write Info files, to check that they are outputting the correct locations. Due to bugs and malfeasances in the various Info writing programs over the years and versions, it is not advisable to ever use this option when just trying to read documentation.

--subnodes

This option only has meaning when given in conjunction with --output. It means to recursively output the nodes appearing in the menus of each node being output. Menu items which resolve to external Info files are not output, and neither are menu items which are members of an index. Each node is only output once.

-v name=value
--variable=name=value

Set the info variable name to value. See Chapter 13 [Variables], page 27.

--version

Prints the version information of Info and exits.

--vi-keys

This option binds functions to keys differently, to emulate the key bindings of vi and Less. The bindings activated by this option are documented in Section 15.1 [infokey format], page 33. (See Chapter 15 [Custom Key Bindings], page 33, for a more general way of altering GNU Info’s key bindings.)

--where
--location

-w Show the filename that would be read and exit, instead of actually reading it and starting Info.

Finally, Info defines many default key bindings and variables. See Chapter 15 [Custom Key Bindings], page 33, for information on how to customize these settings.

13 Manipulating Variables

GNU Info uses several internal variables whose values are looked at by various Info commands. You can change the values of these variables, and thus change the behavior of Info, if desired.

There are three ways to set the value of a variable, listed here in order of precedence:

1. interactively, using the set-variable command described below;

2. on the command line, using the -v (--variable) command line option (see [variableassignment], page 25);

3. in the #var section of the .infokey file (see Chapter 15 [Custom Key Bindings], page 33).

M-x set-variable

Read the name of a variable, and the value for it, in the echo area and then set the variable to that value. Completion is available when reading the variable name (see Section 9.3 [The Echo Area], page 14); completion is also available when reading the value when that makes sense.

M-x describe-variable

Read the name of a variable in the echo area and display its value and a brief description.

Here is a list of the variables that you can set in Info.

automatic-footnotes

When set to On, footnotes appear and disappear automatically; else, they appear at the bottom of the node text. This variable is Off by default. When a node is selected, a window containing the footnotes which appear in that node is created, and the footnotes are displayed within the new window. The window that Info creates to contain the footnotes is called *Footnotes*. If a node is selected which contains no footnotes, and a *Footnotes* window is on the screen, the *Footnotes* window is deleted. Footnote windows created in this fashion are not automatically tiled so that they can use as little of the display as is possible.

automatic-tiling

When set to On, creating or deleting a window resizes other windows. This variable is Off by default. Normally, typing ‘C-x 2’ divides the current window into two equal parts. When automatic-tiling is set to On, all of the windows are resized automatically, keeping an equal number of lines visible in each window. Any *Completions* and *Footnotes* windows are exceptions to the automatic tiling; they retain their original size.

cursor-movement-scrolls

When set to On, when cursor movement commands reach the top or bottom of a node, another node is loaded depending on the value of scroll-behavior (see below). This is the default. When this variable is set to Off, cursor movements stop at the top or bottom of a node.

errors-ring-bell

When set to On (the default), errors cause the bell to ring.

follow-strategy

When set to remain (the default), Info tries to remain within the directory containing the currently displayed Info file when following a cross-reference to an external manual, before looking for the referenced manual in the search path. The alternative value is path, which means to look through the search path right away. remain is intended to be useful for several Texinfo manuals that all reference each other and whose versions should match each other. (For example, various manuals relating to a particular version of Emacs.) The alternative behavior, with path, may be useful when your Info file search path parallels your command shell’s search path, and you always want to find documentation of the version of the program that the shell would execute.

gc-compressed-files

When set to On, Info garbage collects files which had to be uncompressed. The default value of this variable is Off. Whenever a node is visited in Info, the Info file containing that node is read into memory, and Info reads information about the tags and nodes contained in that file. Once the tags information is read by Info, it is never forgotten. However, the actual text of the nodes does not need to be retained unless a particular Info window needs it. For non-compressed files, node text is not remembered when it is no longer in use. But de-compressing a file can be a time-consuming operation, and so Info tries hard not to do it twice. This variable tells Info it is okay to garbage collect the text of the nodes of a file which was compressed on disk.

hide-note-references

By default, Info displays the contents of Info files mostly verbatim, including text that is used by Info readers for navigation (for example, marking the location of menus or cross-references). If you set this variable to On, some of this text is hidden, in a similar way to the Info-hide-note-references variable in Emacs (see Section “Emacs Info Variables” in Info).

highlight-searches

When set to On, highlight matches from searching commands (see Chapter 6 [Searching Commands], page 8).

infopath-no-defaults

Used in conjunction with the INFOPATH environment variable (see [INFOPATH], page 23). When set to On, the default documentation directory defined when Info was built (e.g., /usr/share/info) is not added to the search path for Info files.

ISO-Latin

The default is On, which means that Info accepts and displays characters represented by bytes with values 128 and above, such as characters in the UTF-8 encoding or in various 8-bit ISO Latin characters, as well as allowing you to input such characters.

The only reason to set this variable to Off would be if your terminal set the eighth bit of a byte to represent the Meta key being pressed.

key-time Length of time in milliseconds to wait for the next byte of a byte sequence generated by a key (or key chord) on the keyboard. For example, if the down key generates the byte sequence ESC O B, and the two bytes ESC O have been received, then a B byte would have to be received within this length of time for a key press of down to be registered. You may wish to set this variable to a larger value for slow terminals or network connections. If you set this variable to 0, it’s unspecified whether a recognized byte sequence representing a key takes precedence over another recognized sequence representing a key that is an initial subsequence of the first sequence. In some cases, you may be able to make pressing a special key on the keyboard that Info doesn’t know about (for example, a function key) cause a command to be executed by setting this variable to 0, and giving the byte sequence the key sends in .infokey. (See Chapter 15 [Custom Key Bindings], page 33.)

min-search-length

Minimum length of a search string (default 1). Attempts to initiate a search for a string (or regular expression) shorter than this value, result in an error.

mouse What method to use to get input from a mouse device. The default value is ‘Off’. Set this variable to normal-tracking to make Info use “normal tracking mode” if it detects that the terminal supports it. This enables you to scroll the contents of the active window with a mouse scrollwheel. On terminal emulators running under the X Window System, such as xterm, you can usually select text with the mouse. However, mouse tracking mode may interfere with this. When this happens, you may be able to select text by holding down the Shift key while clicking and dragging.

nodeline How to print the node header line that appears at the top of each node. By default only the pointers to neighbouring nodes are displayed (the “Next”, “Prev”, and “Up” pointers): this corresponds to the pointers value for this variable. To print the entire line, set nodeline to the value print, which will include the filename and name of the node. To not display the header line at all, use the value no.

scroll-behavior
scroll-behaviour

The two variable names are synonymous. Control what happens when scrolling commands are used at the end or beginning of a node (see Chapter 5 [Scrolling Commands], page 6). The default value for this variable is Continuous. Possible values:

Continuous

Try to get the first item in this node’s menu, or failing that, the ‘Next’ node, or failing that, the ‘Next’ of the ‘Up’ node. This behavior is identical to using the ‘]’ (global-next-node) and ‘[’

          (global-prev-node) commands.

Next Only Only try to get the ‘Next’ node.

Page Only Just stop, changing nothing. With this value, no scrolling command can change the node that is being viewed.

This variable also affects cursor movement commands (see Chapter 4 [Cursor Commands], page 5) unless the cursor-movement-scrolls variable is set to Off. See [cursor-movement-scrolls], page 27.

scroll-last-node

Control what happens when a scrolling command is issued at the end of the last node. Possible values are:

Stop  Do not scroll. Display the ‘No more nodes within this document’

message. This is the default.

Top Go to the top node of the document.

This variable is in effect only if scroll-behavior is set to Continuous.

scroll-step

The number of lines to scroll to bring the cursor back into the window. The default value of this variable is 1, which causes a kind of “smooth scrolling” which some people prefer. Scrolling happens automatically if the cursor has moved out of the visible portion of the node text. If the variable scroll-step is 0, the cursor (and the text it is attached to) is placed in the centre of the window.

search-skip-screen

Set the starting point of repeated searches (see [repeated-search], page 8). When set to Off (the default), repeated searches start at the position immediately following (when searching in forward direction), or immediately preceding (when searching backwards) the cursor. When set to On, repeated searches omit lines visibly displayed on the screen. In other words, forward searches (}) start at the beginning of the next page, and backward searches ({) start at the end of the previous page.

show-index-match

When set to On (the default), the portion of the matched search string that you typed is indicated (by displaying it in the “opposite” case) in the result message (see Chapter 6 [next-index-match], page 8).

visible-bell

When set to On, Info attempts to flash the screen instead of ringing the bell. This variable is Off by default. If the terminal does not allow flashing, this variable has no effect. (But you can still make Info perform quietly by setting the errors-ring-bell variable to Off; or using an external command to mute the bell, e.g., xset b 0 0 0.)

14 Colors and Styles

You can choose to highlight parts of Info’s display, such as cross-references and search matches, using a variety of styles, including colors, boldface and underline. Here are the variables that are available to do this:

link-style

Used for cross-references and menu entries.

active-link-style

Used for a cross-reference or menu entry when typing RET would have the effect of following said cross-reference or menu entry.

match-style

Used for matches from a search command. (See Chapter 6 [Searching Commands], page 8.)

Each of these is given in the .infokey file just as the variables in the previous chapter. Their values are a comma-separated list of values in the following table:

black
red
green
yellow
blue
magenta
cyan

white Use the color specified for text.

nocolor

nocolour Turn off any color that was in effect, using the terminal’s default color.

bgblack
bgred
bggreen
bgyellow
bgblue
bgmagenta
bgcyan

bgwhite Use the color specified for the background.

bgnocolor
bgnocolour

Use the terminal’s default background color.

underline
nounderline

Turn text underline on or off.

standout
nostandout

Turn ‘standout mode’ on or off. Standout mode entails the use of appearance modes that make text stand out, and varies between terminals.

bold
regular

nobold Turn boldface on or off.

blink

noblink Make the text blink, or not.

Here is an sample excerpt from an .infokey file:

     #var
     link-style=yellow
     active-link-style=yellow,bold
     match-style=underline,bold,nocolor

With this, cross-references are all yellow, and active cross-references are additionally displayed in bold. Any search matches will be shown in bold, and underlined. Moreover, if there is a search match inside a cross-reference, the ‘nocolor’ rendition style will cancel the yellow color, leaving the text in the match the terminal’s default color. (Note, however, that the rendition styles for active cross-references take priority over those for search matches, so search matches there will still be displayed in yellow.)

15 Customizing Key Bindings and Variables

Info allows you to override the default key-to-command bindings and variable settings described in this document. (The --vi-keys option rebinds many keys at once; see [–vi-keys], page 26.)

On startup, GNU Info looks for a configuration file in the invoker’s HOME directory called .infokey, i.e., ~/.infokey.^{1} If it is present, then Info adopts the key bindings and variable settings contained therein. To use an alternative configuration file, use the --init-file option (see [–init-file], page 24).

Variables may also be set on the command line with the --variable option (see [variableassignment], page 25). Variable settings on the command line override settings from the

.infokey file.

15.1 infokey format

Here is an example .infokey file which specifies the key bindings that are activated by the --vi-keys option to Info (see [–vi-keys], page 26).

#info
g       first-node
G       last-node
\mb     beginning-of-node
\me     end-of-node
j       next-line
k       prev-line
f       scroll-forward-page-only
^f      scroll-forward-page-only
\m\     scroll-forward-page-only
z       scroll-forward-page-only-set-window
b       scroll-backward-page-only
^b      scroll-backward-page-only
w       scroll-backward-page-only-set-window
\kd     down-line
^e      down-line
^j      down-line
^m      down-line
\ku     up-line
^y      up-line
^k      up-line
d       scroll-half-screen-down
^d      scroll-half-screen-down

1 Due to the limitations of DOS filesystems, the MS-DOS version of Info looks for a file _infokey instead.

If the HOME variable is not defined, Info additionally looks in the current directory.

u       scroll-half-screen-up
^u      scroll-half-screen-up
^xn     next-node
^xp     prev-node
^xu     up-node
'       last-node
\mt     top-node
\md     dir-node
^xg     goto-node
I       goto-invocation-node
n       search-next
N       search-previous
\mf     xref-item
^xr     xref-item
\mg     select-reference-this-line
^x^j    select-reference-this-line
^x^m    select-reference-this-line
^c      abort-key
\mh     get-info-help-node
:q      quit
ZZ      quit
#echo-area
\mh     echo-area-backward
\ml     echo-area-forward
\m0     echo-area-beg-of-line
\m$     echo-area-end-of-line
\mw     echo-area-forward-word
\mx     echo-area-delete
\mu     echo-area-abort
^v      echo-area-quoted-insert
\mX     echo-area-kill-word

The file consists of one or more sections. Each section starts with a line that identifies the type of section. The possible sections are:

#info Key bindings for Info windows. The start of this section is indicated by a line containing just #info by itself. If this is the first section in the source file, the #info line can be omitted. The rest of this section consists of lines of the form:

          string whitespace action [ whitespace [ # comment ] ] newline

Whitespace is any sequence of one or more spaces and/or tabs. Comment is any sequence of any characters, excluding newline. string is the key sequence which invokes the action. action is the name of an Info command. The characters in string are interpreted literally or prefixed by a caret (^) to indicate a control character. A backslash followed by certain characters specifies input keystrokes as follows: \b Backspace \e Escape (ESC) \n Newline \r Return \t Tab \ku Up arrow \kd Down arrow \kl Left arrow \kr Right arrow \kU Page Up \kD Page Down \kh HOME

\ke  END

\kx Delete (DEL) \mx Meta-x where x is any character as described above. Backslash followed by any other character indicates that character is to be taken literally. Characters which must be preceded by a backslash include caret, space, tab, and backslash itself.

#echo-area

Key bindings for the echo area. The start of this section is indicated by a line containing just #echo-area by itself. The rest of this section has a syntax identical to that for the key definitions for the Info area, described above.

#var Variable initializations. The start of this section is indicated by a line containing just #var by itself. Following this line is a list of variable assignments, one per line. Each line consists of a variable name (see Chapter 13 [Variables], page 27) followed by = followed by a value. There may be no white space between the variable name and the =, and all characters following the =, including white space, are included in the value.

Blank lines and lines starting with # are ignored, except for the special section header lines.

Key bindings defined in the .infokey file take precedence over GNU Info’s default key bindings, whether or not ‘--vi-keys’ is used. A default key binding may be disabled by overriding it in the .infokey file with the action invalid. In addition, all default key bindings can be disabled by adding this line anywhere in the relevant section:

     #stop

This will cause GNU Info to ignore all the default key commands for that section.

Beware: #stop can be dangerous. Since it disables all default key bindings, you must supply enough new key bindings to enable all necessary actions. Failure to bind any key to the quit command, for example, can lead to frustration.

Note that some keys (such as C-c) have special meanings to terminals, and any bindings for these would not be effective. See Section “Special Characters” in GNU Coreutils.

The order in which key bindings are defined in the .infokey file is not important, except that the command summary produced by the get-help-window command only displays the first key that is bound to each command.

16 Alternate MS-DOS/Windows keybindings

Here are some keybindings that are only used in Info on MS-DOS/Windows.

Cursor commands

Home      beginning-of-line
End       end-of-line
C-RIGHT   forward-word
C-LEFT    backward-word
C-Home    beginning-of-node
C-End     end-of-node

Node commands

C-NEXT    next-node
C-PREVIOUS
          prev-node
C-UP (an arrow key)
          up-node
C-CENTER  history-node

The NEXT key is known as the PgDn key on some keyboards. The PREVIOUS key is known as the PgUp key on some keyboards.

Echo area commands

C-RIGHT   echo-area-forward-word
C-LEFT    echo-area-backward-word
Shift-TAB
          echo-area-tab-insert

On DOS/Windows only, the Shift-TAB key is an alias for M-TAB. This key is sometimes called ‘BackTab’.

Miscellaneous commands

F1        get-help-window

Appendix A Index

*

*Footnotes* window. . . . . . . . . . . . . . . . . . . . . . . . . . . 27

,

, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

–

- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 --all (-a) command line option . . . . . . . . . . . . . . . 22 --apropos (-k) command line option . . . . . . . . . . . 22 --debug (-x) command line option . . . . . . . . . . . . . 23 --directory (-d) command line option . . . . . . . . 23 --dribble command line option. . . . . . . . . . . . . . . . 23 --file (-f) command line option . . . . . . . . . . . . . . 23 --help (-h) command line option . . . . . . . . . . . . . . 24 --index-search command line option . . . . . . . . . . 24 --init-file command line option . . . . . . . . . . . . . 24 --node (-n) command line option . . . . . . . . . . . . . . 24 --output (-o) command line option . . . . . . . . . . . . 24 --raw-escapes (-R) command line option . . . . . . 24 --restore command line option. . . . . . . . . . . . . . . . 25

--show-malformed-multibytes command line

option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

--show-options (--usage, -O) command line

option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --speech-friendly (-b) command line option. . 25

--strict-node-location command line option. 25

--subnodes, command line option . . . . . . . . . . . . . . 25 --variable (-v) command line option. . . . . . . . . . 25 --version command line option. . . . . . . . . . . . . . . . 26 --vi-keys command line option. . . . . . . . . . . . . . . . 26 --where (--location, -w) command line option. 26

.

.infokey format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

/

/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

=

=, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

?

? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 ?, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

[

[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

]

] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

_info file (MS-DOS) . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

{

{ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

}

} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

0

                                                0 ... 9, vi-like operation . . . . . . . . . . . . . . . . . . . 18

0, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

1

1 … 9, in Info windows . . . . . . . . . . . . . . . . . . . . . . . 11

A

abort-key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 absolute Info file names . . . . . . . . . . . . . . . . . . . . . . . . 21 active-link-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

                                                add-digit-to-numeric-arg . . . . . . . . . . . . . . . . . . . . 18

ANSI escape sequences in documents . . . . . . . . . . . 24 Apropos, in Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 arguments, command line . . . . . . . . . . . . . . . . . . . . . . 21 arguments, negative. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 automatic-footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . 27 automatic-tiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

Appendix A: Index 39

B

b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 BackTab, in Info windows . . . . . . . . . . . . . . . . . . . . . 12 BackTab, in the echo area . . . . . . . . . . . . . . . . . . . . 37 backward-char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 backward-word. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 beginning-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 beginning-of-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Blinking text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Bold text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 BS (backspace) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 bugs, reporting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

C

C-a, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-a, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-b, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-CENTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-d, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-e, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-e, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-End. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-f, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-f, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-g, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-g, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 C-Home. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-k, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-LEFT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-LEFT, in the echo area . . . . . . . . . . . . . . . . . . . . . 37 C-n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-NEXT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-PgDn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-PgUp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-PREVIOUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-q, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 C-RIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-RIGHT, in the echo area . . . . . . . . . . . . . . . . . . . . 37 C-s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-t, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 C-u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 C-UP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 C-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 C-w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x ^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-x 0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-x 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-x 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-x b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-x C-b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-x C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 C-x C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-x DEL, in the echo area. . . . . . . . . . . . . . . . . . . . . 15 C-x n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-x N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 C-x o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 C-x t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 C-y, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 16 cancelling the current operation . . . . . . . . . . . . . . . . 19 cancelling typeahead . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 case-sensitive search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 case-sensitivity, and search . . . . . . . . . . . . . . . . . . . . . . 9 clear-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Colored background. . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Colored foreground . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 colors in documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 command line options . . . . . . . . . . . . . . . . . . . . . . . . . . 21 command-line options, how to find . . . . . . . . . . . . . 25 commands, describing . . . . . . . . . . . . . . . . . . . . . . . . . . 18 completion. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 compressed Info files . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 current file, information about . . . . . . . . . . . . . . . . . 18 cursor, moving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

                                                cursor-movement-scrolls . . . . . . . . . . . . . . . . . . . . . 27

customizing key bindings. . . . . . . . . . . . . . . . . . . . . . . 33

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 default key bindings, overriding . . . . . . . . . . . . . . . . 33 DEL, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 DEL, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 delete-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 describe-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 describe-key. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 describe-variable . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 dir-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 directory path . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 display-file-info . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 down-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 DOWN (an arrow key) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

E

e, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 echo area . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 echo-area-abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 echo-area-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

                                                echo-area-backward-kill-line . . . . . . . . . . . . . . . 15
                                                echo-area-backward-kill-word . . . . . . . . . . . . . . . 15
                                                echo-area-backward-word . . . . . . . . . . . . . . . . . . . . . 15

echo-area-beg-of-line . . . . . . . . . . . . . . . . . . . . . . . 14 echo-area-complete . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-delete . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 echo-area-end-of-line . . . . . . . . . . . . . . . . . . . . . . . 15 echo-area-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 echo-area-forward-word . . . . . . . . . . . . . . . . . . . . . . 15 echo-area-kill-line. . . . . . . . . . . . . . . . . . . . . . . . . . 15 echo-area-kill-word. . . . . . . . . . . . . . . . . . . . . . . . . . 15 echo-area-newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

echo-area-quoted-insert . . . . . . . . . . . . . . . . . . . . . 15

echo-area-rubout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

echo-area-scroll-completions-window . . . . . . . 16

echo-area-tab-insert . . . . . . . . . . . . . . . . . . . . . . . . 15

echo-area-transpose-chars. . . . . . . . . . . . . . . . . . . 15

echo-area-yank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 echo-area-yank-pop . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Emacs Info reader. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 End . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 end-of-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 end-of-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 errors-ring-bell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ESC C-f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 ESC C-v, in Info windows . . . . . . . . . . . . . . . . . . . . . 14 ESC C-v, in the echo area. . . . . . . . . . . . . . . . . . . . . 16 execute-command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

F

f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 F1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 file names, relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 file, outputting to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 files, compressed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 find-menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 finding the Invocation node . . . . . . . . . . . . . . . . . . . . . 3 first-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 follow-strategy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 footnotes window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 footnotes, displaying . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 format of .infokey . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 forward-char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 forward-word . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 functions, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

G

g . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 gc-compressed-files. . . . . . . . . . . . . . . . . . . . . . . . . . 28 get-help-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 get-info-help-node . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 global-next-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 global-prev-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 goto-invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 goto-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 grow-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 hide-note-references . . . . . . . . . . . . . . . . . . . . . . . . 28 highlight-searches . . . . . . . . . . . . . . . . . . . . . . . . 9, 28 history-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Home . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

I

i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 incremental search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 index search, selecting from the command line . . 24 index, searching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 index, virtual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 index-apropos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 index-search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 indexes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 indices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Info files, compressed . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Info files, reading in Emacs . . . . . . . . . . . . . . . . . . . . . . 1 Info files, relative . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Info files, searching all indices . . . . . . . . . . . . . . . . . . 22 Info manual location . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Info manual, specifying initial . . . . . . . . . . . . . . . . . . 23 Info, invoking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 info-version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 INFO_PRINT_COMMAND, environment variable . . . . . 17 infokey format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 infokey, program for customizing key bindings . 33 infopath-no-defaults . . . . . . . . . . . . . . . . . . . . . . . . 28 INFOPATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 initial node, specifying . . . . . . . . . . . . . . . . . . . . . . . . . 23 invocation description, how to find . . . . . . . . . . . . . 25 invoking Info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 isearch-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 isearch-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ISO Latin characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 ISO-Latin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

K

keep-one-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 key bindings, customizing . . . . . . . . . . . . . . . . . . . . . . 33 key-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 keys, describing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 keystrokes, recording . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

L

l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 last-menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 last-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 LEFT (an arrow key) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LEFT, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . 14 Less-like key bindings . . . . . . . . . . . . . . . . . . . . . . . . . . 26 link-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 list-visited-nodes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 local printer device. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 M-- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 M-< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-0 ... M-9. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 M-0, vi-like operation . . . . . . . . . . . . . . . . . . . . . . . 12 M-1 … M-9, vi-like operation . . . . . . . . . . . . . . . . . 11 M-b, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-b, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 M-BS, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . 15 M-d, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 M-DEL, in the echo area. . . . . . . . . . . . . . . . . . . . . . . 15 M-f, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-f, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 M-r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 M-TAB, in Info windows . . . . . . . . . . . . . . . . . . . . . . . 12 M-TAB, in the echo area. . . . . . . . . . . . . . . . . . . . . . . 15 M-v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 M-y, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 16 malformed multibyte sequences, showing . . . . . . . 25 man . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 man pages, displaying . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 match-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 menu, following . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 menu, following, from inside Info . . . . . . . . . . . . . . . . 4 menu-digit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 menu-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 menu-sequence. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Meta key sets eighth bit. . . . . . . . . . . . . . . . . . . . . . . . 28 min-search-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 mouse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 move-to-next-xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 move-to-prev-xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 move-to-window-line. . . . . . . . . . . . . . . . . . . . . . . . . . . 5 moving the cursor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 negative arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 next-index-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 next-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 next-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 next-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 NEXT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 node header line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 node, selecting from the command line . . . . . . . . . 24 nodeline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 nodes, selection of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Non-interactive usage . . . . . . . . . . . . . . . . . . . . . . . . . . 22 numeric arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 numeric arguments, negative . . . . . . . . . . . . . . . . . . . 19

O

online help, using Info as . . . . . . . . . . . . . . . . . . . . . . . 24 options, command line . . . . . . . . . . . . . . . . . . . . . . . . . 21 outputting to a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 overriding default key bindings . . . . . . . . . . . . . . . . . 33 O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 PageDown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 PageUp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 prev-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 prev-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 prev-window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 PREVIOUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 print-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 printing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

                                                printing characters, in the echo area . . . . . . . 15

printing nodes to the local printer . . . . . . . . . . . . . . 17

Q

q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 quitting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

R

r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 redraw-display . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 regular expression search . . . . . . . . . . . . . . . . . . . . . . . . 8 relative Info file names . . . . . . . . . . . . . . . . . . . . . . . . . 21 remembering user keystrokes . . . . . . . . . . . . . . . . . . . 23 repeated search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 replaying recorded keystrokes . . . . . . . . . . . . . . . . . . 25 R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 RET, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . 12 RET, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 15 RIGHT (an arrow key) . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 RIGHT, in the echo area. . . . . . . . . . . . . . . . . . . . . . . 14

S

s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 screen, changing the height of . . . . . . . . . . . . . . . . . . 19 scroll-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

                                                scroll-backward-page-only . . . . . . . . . . . . . . . . . . . . 6
                                                scroll-backward-page-only-set-window . . . . . . . 6

scroll-behavior . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 scroll-behaviour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 scroll-forward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

                                                scroll-forward-page-only . . . . . . . . . . . . . . . . . . . . . 6
                                                scroll-forward-page-only-set-window . . . . . . . . 6
                                                scroll-half-screen-down . . . . . . . . . . . . . . . . . . . . . . 7

scroll-half-screen-up . . . . . . . . . . . . . . . . . . . . . . . . 7 scroll-last-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 scroll-other-window. . . . . . . . . . . . . . . . . . . . . . . . . . 14 scroll-step . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 scrolling through node structure . . . . . . . . . . . . . . . . . 7 search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 search, and case-sensitivity . . . . . . . . . . . . . . . . . . . . . . 9 search, case-sensitive . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 search-backward . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

search-case-sensitively . . . . . . . . . . . . . . . . . . . . . . 8

search-next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 search-previous . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 search-skip-screen . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Searching all indices . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 searching, in the indices . . . . . . . . . . . . . . . . . . . . . . . . 10

select-reference-this-line. . . . . . . . . . . . . . . . . . 12

select-visited-node. . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Selecting text with the mouse . . . . . . . . . . . . . . . . . . 29 set-screen-height . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 set-variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

Shift-TAB, in Info windows . . . . . . . . . . . . . . . . . . . 12
Shift-TAB, in the echo area . . . . . . . . . . . . . . . . . . 37

show-footnotes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 show-index-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 slow network connections . . . . . . . . . . . . . . . . . . . . . . 29 SPC, in Info windows . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 speech synthesizers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 split-window. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 Standard output not a terminal . . . . . . . . . . . . . . . . 22 startup node, specifying . . . . . . . . . . . . . . . . . . . . . . . . 23 S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

T

t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 TAB, in Info windows. . . . . . . . . . . . . . . . . . . . . . . . . . 12 TAB, in the echo area . . . . . . . . . . . . . . . . . . . . . . . . . 16 tile-windows. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 tiling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 toggle-regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 toggle-wrap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 top-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 tree-search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 tree-search-next . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 tree-search-previous. . . . . . . . . . . . . . . . . . . . . . . . . . 9

U

u . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Underlined text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 universal-argument . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 up-line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 up-node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 UP (an arrow key) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

V

variable assignment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 variables, describing. . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 variables, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 version information . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 vi-like key bindings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 view-file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 virtual-index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 visible-bell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

W

Where is an Info manual? . . . . . . . . . . . . . . . . . . . . . . 26 where-is . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 windows, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 windows, deleting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 windows, manipulating . . . . . . . . . . . . . . . . . . . . . . . . . 13 windows, selecting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

X

xref-item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 xterm mouse selections. . . . . . . . . . . . . . . . . . . . . . . . . 29